An advanced computer sends information to its various parts via infrared light pulses traveling...

If the information has to be obtained from memory in 2.0 ns, what is the maximum distance the memory unit can be from the central processing unit?
Answer #1
Given that the tiem taken is t = 2.0 ns = 2.0*10^-9s
The speed of signal c = 3.0*10^8 m/s
(Speed of light)
The speed of light in silicon isv = c / nwhere n = 4.01 isthe refractive index of silicon.v = 3.0*10^8 m/s / 4.01= 0.748*10^8 m/s
Then the distance is d = c*t
d = (0.748*10^8 m/s)(2.0*10^-9s )
=0.1496 m
answered by: Roman